February FEL Monthly Highlights


We successfully operated the IR Demo the entire month for the winter FEL user run.  Useful data was obtained by all the FEL users (see the “Operations” section below for details) and for our own measurements of system performance for the Upgrade.

With regard to Upgrade system tests, we successfully ran all eight (8) klystrons in Zone 4 at 5 kW simultaneously.

We prepared for the semiannual review of the Upgrade project scheduled for March 27-28.


We also completed a draft of a new cost performance plan that will accommodate the accumulated ECR’s during the current Phase 1 of the project and the new Phase 2 of the project commencing this month.  We had this draft reviewed by the contract monitor so that it can be used for the project cost performance reports starting with this February report.

The Upgrade Program Review Committee has chosen March 27 and 28 for the next semi-annual review of the Upgrade Program.  The first day of the review will be a status report by each WBS Manager.  The second day of review will be a workshop on scaling issues that will include contributions from the members of the review panel.

F. Dylla and C. Leemann had the pleasure of meeting the Chief of Naval Research, Admiral Jay Cohen at the Feb. 5-6 meeting of the Virginia Research and Technology Commission in Richmond.  During off line
discussions Adm. Cohen expressed his interest in the FEL program and his desire to visit the lab in the near future for a tour and briefing.  We submitted on Feb. 5, jointly with VCU, a $3.75 M proposal to the
Commonwealth Technology and Research Fund for recommissioning the x-ray synchrotron donated by one of the FEL industrial partners.  Also three proposals went to NSF during February by FEL users to support FEL user lab infrastructure.

On Feb. 15 we participated in a workshop sponsored by the NSF and the ODU School of Engineering.  ODU is a finalist in the competition for an Industry University Cooperative Research Center (IUCRC) to be hosted in the Applied Research Center.  The proposed center would focus on laser processing.

Project Cost Performance:

The project (Phase 1) budget for the period June 1, 2000 to Sept. 30, 2001 is $9,029k.  The Phase 2 project was approved for $4,500 k for a performance period of Feb. 1, 2001 to Sept. 30, 2002.  With the approval of the Phase 2 project, a rebaselined cost performance plan was developed and approved by the ONR contract monitor.  Cost performance will now be reported against both the Phase 1 plan and the Phase 2 plan when Phase 2 charges are accumulated.  For the convenience of reviewers, a roll-up of total project performance for both Phase 1 and Phase 2 will be presented.  The project through the month of February has a total of $2,044k of performance scheduled (assuming the project started at the
originally planned start date of April 1, 2000).  The work performed through the end of February was $2,998k, which is 33% complete vs. 23% scheduled.  The actual cost accrued through February totals $3,425k.  This results in a schedule variance of +$953k and a cost variance of -$427k.

Note:  This rollup is different from the cost performance summary shown at the March 27 review, which included an error in earned value (BCWP) claimed for the RF systems (WBS 6) noted in the review and now corrected.  The errant earned value will be claimed at a later date when the RF hardware is delivered.  There are no changes in the budgeted cost of work scheduled (BCWS) or the actual cost (ACWP).

WBS 3 (Beam Physics):

B. Yunn evolved an injector solution based on the 500 keV Engwall gun, sporting very low emittance (~2;3.5 mm-mrad/16;32 pi deg-keV) at high charge state (~135;270 pC).  This is adequate for the UV driver and likely will support JERBIL (D. Douglas's concept for a 4th generation light source).  Work continued on interfacing the UV bypass, with most effort focussed on management of momentum compaction while avoiding impressively strong trim quads.

A linac to linac UV bypass solution was completed by the end of the month.  It provides for placement of the wiggler in the pit.  Preliminary analysis indicates single particle performance (chromatic, geometric) is adequate, but analysis is ongoing.  Machine error, CSR, and other collective effects have not yet been investigated.

WBS 4 (Injector):

We started planning for gun testing at the end of the 2001.

We continued work on field emission reduction.  The FET system will be modified to collect more data.  Refinement of the internal cesiation scheme was started.  The first cut of the gun chamber drawings were
completed this month.

Gun HVPS - Resolved the differences with the vendor concerning the multiplier stacks.  There is just no other alternative to 2 multiplier stacks and 1 monitor stack.  This means the HVPS tank will have to be
extended for the additional multiplier stack.

WBS 5 (SRF):

Two (2) cavities fabricated this month; this makes four (4) complete. First article helium vessels are undergoing inspection and QA.  HOM component design completed and components are being fabricated.  Stress test of the beamline flange completed.  Flange was deflected 50 mrad and
remained leak tight at 2K.

Prototype tuner #2 test and evaluation is underway.  Received the deep draw fixture for the fundamental power coupler body and QA looks good.  Question back out to thermal shield vendors from initial proposals. Detailing cold valve details with vendor for quote.

WBS 6 (RF):

Zone 3
Waiting on part delivery before further assembly and test progress on this system.
Zone 4:
Work continues on the test software to complete and document the zone.  All 8 klystrons have been run at 5 kW simultaneously.  The calibration tests have not yet been run due to incomplete calibration software.  This software is in progress.  A “punch list” of 12 items needs to be completed before the zone is finished.
Injector RF
- Ordered 2 autotransformers as backups for the existing 3 phase Variacs.  PSC is still considering the cost and schedule impact of using autotransformers in the upgrade.  Reliability and space
requirements should improve.
-CPI is having a problem getting their vendors deliver the various parts.  It appears the first 100 kW klystron delivery will be about a month late at 5/18/01.  This will have no impact on the program.

WBS 8 (I&C):

Two new optical BPM detector assemblies were built and tested and are available as well as new cables for testing the BPM system  Two new local shutter control boxes for user labs are complete and await
testing.  Also in progress is the construction of power supply racks for each of the user labs. Drawings are in to EECAD for GPIB System drawing and an update to the LSS User Interface Box in the user labs.

The first article (2 pcs) of the 3" shielded beam viewer are in an being tested; they look great and have passed leak detection.  The residual gas analyzer (RGA) output showed no contaminates.  The 3" BPMs are due in a week.

A divide-by-two was added to the 37 MHz monitor in the clean room to produce a clean 18 MHz reference synchronous with the FEL micropulses, thanks to Hovator, et al.

Discussions continue for commercial trim cards.  A number of companies have expressed interest.  The motivation for changing is a higher current is needed and the replacement cost for the 10 amp cards
(complete 32 channel rack) is ~ $4000 per channel - my target price is <$2000.  The new system would also address real time monitoring - the current system can not tell if the current had fluctuated putting you
off loop.

The new picomotor controller chassis, which will control the transport mirrors, has been tested on the bench and is ready for installation.  The parts continue to arrive for the vacuum control system and power
supplies. Fabrication shops are being queried for the ion pump power supplies assembly.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o  DULY Research continues to model the Small Injector dipole.  They are working the on the 3 D
    model as well as the 2D details.
o  DULY Research resolved their magnetic model’s anomalous tapered field results by going to a finer
    mesh.  They continue working the on the 3D drafting model as well as the 2D details.
o  DULY Research obtained a field that was within the flatness specification using their 3D model  (in
    RADIA) of a parallel faced dipole and mu metal.  Their results with the actual wedge model still
    have a gradient that is an order of magnitude out of specification, and (counter intuitively) still high on
    the narrow side.  They now have enough confidence in their model that they will make changes to
    investigate the source of the gradient.  Their drawing set and the drafting model of the GV (Small
    injector dipole) include poles, field clamps and coils.  Work is ongoing on the leads and water

Optical Chicane Dipole (DW)
o  We checked the drawing package and started incorporating the information.  We started on the
    specifications for making the magnet.
Arc Dipoles (GY, GX, GQ)
o  AES started the layout the coil's leads, incorporating the 4-in-hand winding and started re directing the
    leads so the water connections could be made while incorporating the bus bars and lead stabilization
o  AES modeled the magnetics of the GY using axisymmetric 2D model which makes provision for the
    out-of-plane curvature of the magnet.  They found that the field is concentrated in the just the inner
    return leg such that the flatness of the field warps to several times the allowed value within the nominal
    good field width.  They will now work on a model with increased width of return leg only to see if the
    flatness returns.  Their next line of action is to increase the thickness of the pole plates.
o  We worked out a way for the leads of the Path Length Correctors to be mounted like the leads of the
    IR Demo's injection dipoles.
o  AES continued the axisymmetric 2D magnetic model of the GY, finding there is indeed a 8.5 to 10 inch
    wide good field region over the full range of excitation using silicon steel boundaries that are within the
    coil on the inside edge and in the gap on the outside edge.  They are now creating a 3D magnetic model
    for final magnetic verification.
QX (3.125" Quad)
o The procurement of steel was started.
o At a working session, we settled on a method of machining fiducial surfaces on the quadrupole that
    satisfies the ability to Wire Electric Discharge Machine the pole tips, provides a mount for the magnets
    on the test stand, allows simple direct mounting to single and double girders and provides for use of
    a bubble level and tooling balls during final alignment.  Sometimes committees actually work!
o Design Package: In parallel with his GW design efforts, N. Vaughan is working on finalizing the QX
    drawings: The drawing package is being modified to incorporate the decisions of the working session.
o Prototype Measurement: Effective length measurements were made using a hall probe.
o Manufacturing philosophy: A meeting with Will Oren, head of engineering technicians and shops resulted
    in a decision to have separate procurement of coils and a separate procurement of cores.  Then, once
    assembly drawings are available, we will probably procure assembly from the resulting vendors but
    keep open the option of assembling in-house.
Sextupole with H/V Correctors
o BT Layout: Layouts are in progress to optimize the use of single
and/or double quad girders.
o Robin Wines continued work on the magnetic model of the Sextupole,
finding the automatic mesh generator makes meshes that yield unallowed
harmonics, a clear indication of improper use of the model. She is
working on creating a mesh that has boundaries that yield zero unallowed
o David Douglas clarified the specifications of the sextupole,
significantly tightening them to 1 part per thousand over the good field
region.  We will need the properly made mesh (above) to resolve errors
of this level.
o We reworked the budget such that the 180° bends, the optical chicane bends and the reverse bend
    chambers will be bought with Phase 1 funds.  This will have to increase the level of support from
    engineering for the chambers in order to achieve these critical procurements.
o The layout of the UV add-on was superposed on the overall machine layout and found to match so no
    changes to the IR machine were necessary.
o Girder arrangements are firming up.  We will make individual girders for each quadrupole, re-use the
    old stands and cartridges, and use long beam pipes through the triplets, (forcing alignment to do a rough
    alignment before the tube is inserted).  The arrangements of diagnostics for the IR machine were
    simplified as the layout progressed.  The result was greater commonality of stands and girders leading
    to simplified design.
o We defined the alignment requirements for the dipoles with the help of Chris Curtis.  We will specify
    machined-in tooling ball holes at the factory with possible vendor CMM verification on all magnets
    other than the 180° Bends.  The latter, we will establish the fiducial locations on site.  Injector dipoles
    will have a surface for a machinist’s level.
o  The layout of the IR lattice was extended into the injection region with study of the use of the existing
    dipole girders.

WBS 10 (Wiggler):

A model for the wiggler is being developed in Radia to determine the best solution for the field bias issue. Design is progressing on the measurement stand.  No decision has yet been made (Karn and Benson) on
whether the existing Group-3 Hall Probes are acceptable or whether a dual axis probe needs to be purchased.

The dispersion section coils were received by PECO and assembly to the core has begun.  A Radia model for the optical klystron wiggler has been developed and is being debugged and checked out.  The initial field profile in the center of the wiggler agrees with previous calculations using Poisson.

WBS 11 (Optics):

We finished review of the check prints for the deformable mirror assembly, and released the drawing package.  The design team continues to work on the upgrade optical transport system (OTS) and optical cavity assemblies.  During operations this month, with extended operation at modest laser power (~ 100 W cw), we found that the O-BPM ADC electronics' limited dynamic range made their use problematic.  We are going to work with I&C on this, and in the short term install a larger range of neutral density filters.  As mentioned in the Jan 8-12 weekly brief, we replaced the black-anodized Al mirror holders on the collimator linear stage with unanodized holders.  This was done because it appeared that most of the beam drift we observed came from the collimator, and the black-anodized mirror holders, with their high CTE, were a likely candidate.  Our operational experience this week has shown a greatly reduced drift, compared with our run in November.  So, this lends credence to our specifications in the upgrade OTS that all mounts be either temperature stabilized or made of low CTE materials.

We also supported operations.  Highlights were:
o  We had the first opportunity to see how well the hutch covering the optical diagnostics worked.  It
    contained the nitrogen atmosphere quite well, greatly reducing measurement problems in the 6 micron
    spectral region.
o  We built a test stand for testing the O-BPM detector assemblies that doesn't involve laser sources.
    This makes testing more flexible.  With it, the O-BPMs in several locations were removed, checked,
    balanced, and, where appropriate, reinstalled.
o We also installed neutral density filters with less attenuation in the filter wheels before the detector

The vacuum vessel for the mirror test stand was welded and leak checked.  We ordered new quadrant pyroelectric detectors from Molectron, and will use them to see if we can get more reliable performance over a larger range of displacements

We received a budgetary estimate on the drawing package for the deformable mirror assembly that is in line with previous estimates.  We will now release a requisition and finish getting quotations.  More
parts for the mirror test stand arrived, namely the new gimbal mount from Newport.  This will be used in the mirror test stand, and will be seriously evaluated for its suitability in the upgrade optical cavity.
We received the new quadrant pyroelectric detectors from Molectron, as well as the rebuilt high power laser beam dump.


The FEL was operated the entire month as the major part of our winter user run.  The machine ran reliably and we had very satisfied users.

Highlights include the following:

G. Luepke (CWM) continued his investigations of the H defects in Si.  The team was able to improve the wavelength and pointing stability of the FEL to improve his signal-to-noise a factor 4-5 compared to his
experimental run in November.

We spent several sessions looking at IR Microscopy using an atomic force microscope.  Ed Gillman (NSU/JLab) obtained our first IR images of a surface using this technique.  It is expected to be able to provide sub-micron resolution on surface absorption in the mid-IR.  Initial experiments were at 5.85 microns.  We also did our first experiments on non-linear dielectric absorption and saw good signals at 18.7 MHz and excellent signal to noise ratios.  Those results are presently under analysis.  We provided several runs for carbon nanotube generation for a NASA/College of William and Mary collaboration at around 6 microns and some interesting results operating the laser with 30 millisecond macropulses.  Overall it is not clear that six micron region is optimal for this work, though.  An answer to that awaits analysis of the nanotube product.

Prof. Robert Austin of Princeton University continued his experiments on energy flow in biomolecules using the amide I aborbance at 6 microns.

We had a successful week at materials processing including studies of laser nitriding and laser hardening. Peter Schaaf from the University of Goettingen continued his investigations of laser nitriding.  He appears
to have successfully nitrided Fe, SS, Al, Ti and Si samples using a simple x-y scanner with a flowing nitrogen purge.  We also delivered light for amorphous metal layer production on iron samples for
Greg Kessel from Virginia Power.